# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 2.2.3 Memorization
Beyond the practical strategy of deduplication, a core theoretical issue in data preparation is the intrinsic mechanism of memorization [wei2024memorization]. While often viewed as a privacy risk, memorization is deeply intertwined with the model’s learning and generalization capabilities. Research in this area has evolved from observing exact replication to analyzing complex memory representations, quantification methods, and its fundamental trade-offs with generalization.
@@
The Mechanism of Memorization. Academic discourse first challenged the traditional view that memorization is caused solely by exact sequence duplication in the training data. A foundational study introduced the concept of “Mosaic Memory” [shilov2024mosaic]. This work posits that LLM memorization is not merely verbatim recall, but a more complex process where models can patch together memories by integrating partially overlapping or similar sequences (i.e., fuzzy duplicates) from the training corpus. Building on this, other research redefined memorization from an adversarial perspective, proposing the “Adversarial Compression Ratio” (ACR) [schwarzschild2024rethinking]. The core idea is that a training sequence is considered “memorized” if it can be elicited by a prompt that is significantly shorter than the string itself. This metric provides a practical, adversarial view for assessing data usage compliance and potential privacy violations.
@@
Quantify the Influence of Memorization. Once complex memory forms were defined, the focus shifted to its quantification and prediction. carlini2022quantifying confirm that memorization is more prevalent than previously believed and is likely to get worse as models continue to scale, at least without active mitigation. This scaling behavior was specifically quantified for factual knowledge, with one study proposing “Scaling Laws for Fact Memorization” [lu-etal-2024-scaling]. It found that a model’s fact knowledge capacity exhibits a linear relationship with model size and a negative exponential relationship with training epochs. Beyond model scale, data-side characteristics are also a critical factor. The “Entropy-Memorization Law” [huang2025entropy] was proposed to investigate the inherent difficulty of memorizing data. This law reveals a linear correlation: the data’s entropy is linearly correlated with its memorization score, suggesting that simpler, lower-entropy data is more easily memorized. In terms of predictability, biderman2023emergent have further shown that using a partially trained model to predict memorization is more effective than using a small model.
@@
Memorization Analysis. The ultimate goal of understanding memorization is to differentiate it from generalization and to enable effective control. A key challenge is diagnosing memorization in black-box models. The PEARL [djire2025memorization] framework was introduced as a novel detection method based on a perturbation sensitivity hypothesis. This hypothesis posits that memorized content is more sensitive to input perturbations, whereas generalized knowledge remains robust. Further analysis reveals a clear trade-off between memorization and generalization across different tasks. Wang et al. [wang2025generalization] traced model capabilities back to pretraining data, finding that task dependencies vary significantly. For instance, Factual Question Answering demonstrates the strongest memorization effect, and this effect increases with model size. Conversely, tasks like machine translation and reasoning exhibit greater generalization, tending to produce novel outputs. Based on these theoretical insights, researchers have begun exploring active mitigation strategies. “Memorization Sinks” [pmlr-v267-ghosal25a] is proposed to activate a unique set of “memorization neurons” for each sequence. This mechanism effectively isolates the memorized content, making it easier to remove without compromising general language capabilities, offering a new path to mitigate the negative impacts of memorization.
@@ 2.3.1 Synthetic Data Generation
One of the most exciting and debated frontiers is the use of synthetic data and the potential for a self-improvement loop [villalobos2024position, long2024llms]. Can a model generate new, high-quality data to train its next generation, thereby kicking off a cycle of recursive self-improvement? While numerous works adopt synthetic data to improve model training [patel2024datadreamer, gilardi2023chatgpt, xu2023knowledge, gandhi2024better, liubest], this idea faces significant theoretical hurdles. A key open question is whether such a process would lead to genuine capability gains or result in model collapse, a degenerative process where the model overfits to its own idiosyncrasies, leading to a gradual loss of diversity and accuracy. Developing a theoretical framework to understand and control the dynamics of this loop is a critical area of research.
@@
Recent researches have started to establish theoretical frameworks for the utility of synthetic data. Gan2024TowardsAT proposes a “reverse-bottleneck” framework, which posits that a post-trained model’s generalization error upper bound is negatively correlated with the “information gain” obtained from the generative model. This suggests that so long as the generative model provides sufficient new information, generalization can in principle be improved. Beyond simple augmentation, the nature of the synthetic data is also being explored. For instance, in the domain of mathematical reasoning, setlur2024rl found that while fine-tuning on correct synthetic answers offers modest gains, using reinforcement learning on the model’s incorrect responses can be twice as sample-efficient. This method helps the model identify and unlearn “spurious correlations” (i.e., incorrect intermediate steps that happen to lead to a correct final answer), ultimately scaling the synthetic dataset’s efficiency by eight-fold compared to standard positive-only finetuning.
@@
The primary theoretical hurdle to recursive self-improvement is “Model Collapse”. shumailov2023curse provide a foundational study on this phenomenon, positing that training on generated data leads to an irreversible degenerative process. Beyond this, other works have highlighted the limitations of synthetic data in capturing human nuance. li-etal-2023-synthetic find that the performance gap between real and synthetic data is smallest for low-subjectivity tasks (like news classification) but much larger for high-subjectivity tasks (like humor or sarcasm detection). This suggests LLMs struggle to generate data with sufficient diversity to capture the complexity of subjective language. This is empirically supported by moller-etal-2024-parrot, which finds that models trained on human-labeled data consistently exhibited superior or comparable performance to those trained on synthetically augmented data.
@@
In response to the threat of model collapse, various mitigation strategies have emerged. Multiple studies have empirically and theoretically demonstrated that the training workflow is the critical factor [gerstgrasser2024is, kazdan2024collapse]. A “replace” workflow, which discards old data and trains new models only on synthetic data, does lead to collapse. However, an “accumulate” workflow, where synthetic data is added alongside the original real data, consistently avoids model collapse and keeps models stable. seddik2024how provided a quantitative estimate, concluding that to maintain stability, the amount of synthetic data used must be considerably smaller than the amount of real data in the training mix. This body of work suggests that the value of synthetic data is highly context-dependent: it can improve performance when real data is scarce but may harm it when real data is plentiful.
@@ 2.3.2 Data Contamination
Data contamination, the inadvertent inclusion of benchmark evaluation samples within the pre-training corpus, poses another critical open challenge in the data preparation stage. This issue fundamentally undermines the validity of model evaluations, making it difficult to discern true generalization capabilities from mere memorization of seen answers [deng-etal-2024-unveiling, cheng2025survey, xu2024benchmark]. The theoretical and empirical investigation of contamination can be broadly categorized by its severe impacts and the methods for its detection and mitigation.
@@
Data contamination is a direct threat to reliable model assessment. Studies demonstrate that contamination can drastically skew evaluation scores. pmlr-v267-kocyigit25a find that contamination in a machine translation task could severely inflate model capability. This work also revealed that larger models exhibit higher sensitivity to contamination, not greater robustness. li2024task further enrich the evaluation, they find that a model’s superior performance in apparent zero- or few-shot settings may not stem from genuine generalization but from its exposure to task-related samples during pre-training. The impact on complex reasoning evaluation is particularly stark. huang2024competition test models on novel competition problems released after their training data cut-off. They find a “cliff like decline” in GPT-4’s performance on medium-to-hard problems, strongly suggesting that its high performance on older benchmarks was reliant on memorization rather than genuine algorithmic reasoning. However, contamination is not limited to harmful verbatim copies. palavalli2024taxonomy establish through experiments that “noisy” or approximate forms of contamination (e.g., masking, augmenting, or noising test examples) can boost performance almost as much as seeing clean, in-domain data. Beyond evaluation, the memorization of contaminated data, especially sensitive information, creates significant privacy vulnerabilities. zhu2024privauditor provides a systematic benchmark for assessing these privacy leakage risks, which are exacerbated during model adaptation and fine-tuning.
@@
In response to these severe impacts, researchers have developed various methods for detection, though effective mitigation remains a significant open problem. pmlr-v267-choi25b propose the KDS framework. Instead of simple text matching, it quantifies contamination by measuring the change in the similarity structure of sample embeddings in the model’s representation space before and after fine-tuning. deng2024investigating introduce an innovative detection method called TS-Guessing. This protocol masks an incorrect answer in a multiple-choice question and prompts the model to fill in the blank. Commercial LLMs were able to “guess” the exact missing wrong option with high accuracy, strongly implying they had memorized the full question format. Other approaches use targeted queries to excavate a model’s memory. chang2023speak use “name cloze” queries to identify a wide range of memorized copyrighted books, which in turn contaminated downstream evaluation tasks. Similarly, dodge2021documenting confirm that the C4 dataset [10.5555/3455716.3455856] contains contaminated examples from NLP benchmarks. Contamination is not just a pre-training issue. tao2025detecting address the challenge of detection after RLHF, where optimization erases traditional likelihood signals. The proposed “Self-Critique” method probes for “policy collapse” by comparing the token-level entropy sequences of an initial response and a second, alternative critique response, where high similarity indicates memorization.
@@ 3.1 Fundamental Problems
(1) How to theoretically evaluate the “power” of a model architecture? This problem focuses on the rigorous analysis of a model’s intrinsic properties, specifically its representability: the capacity to solve or approximate given classes of functions. The challenge is to determine the theoretical limits of architectures under realistic constraints such as finite precision, width, and depth. This involves establishing upper bounds on the model size required to realize specific capabilities and lower bounds that characterize the minimum circuit or communication complexity needed to solve computational tasks.
@@
Of course, discussions on these questions often overlap. For example, according to the “no free lunch” principle, there is usually a trade-off between model performance and computational cost, that is, efficient models may come with potential limitations in representational power. Moreover, the theoretical understanding of a model’s architecture is closely related to the observable characteristics it exhibits. We present a visualization of the topics in this stage in fig. 3. In what follows, we start with the core theories and methods relevant to model design.
@@ 3.2.1 Representability of Models
Recently, the community has renewed interest in the expressive capacity of Transformers, especially in terms of universal approximation [yun2019transformers, jiang2024approximation, kajitsuka2023transformers]. yun2019transformers show that for any sequence-to-sequence function, there exists a Transformer that can approximate it, where the number of layers scales exponentially in model dimension $d$ or input sequence length $T$ and the size of each layer is independent of $d$ and $T$. jiang2024approximation derive explicit Jackson-type approximation bounds for Transformers by introducing new complexity measures to construct appropriate approximation spaces and their results show that Transformers approximate efficiently when the temporal dependencies of the target function exhibit a low-rank structure. kajitsuka2023transformers demonstrate that once the feed-forward layer quantizes continuous inputs, even a one-layer, single-head self-attention module becomes a universal approximator for continuous permutation equivariant functions on a compact domain.
@@
In addition, there are also many works that study representability through Turing completeness, asking whether a model can effectively simulate each step of a Turing machine—and thus inherit the full computational power of Turing-computable problems [dehghani2018universal, perez2021attention, wei2022statistically]. perez2021attention prove that Transformers are indeed Turing complete under the assumption of infinite precision. dehghani2018universal show that standard finite-precision Transformers are not Turing complete and the proposed Universal Transformer which combines parallel self-attention with recurrence can overcome this limitation. wei2022statistically demonstrate that Transformers can statistically meaningfully approximate Turing machines running in time $O(T)$, with sample complexity polynomial in the alphabet size, state-space size, and $\log(T)$. In fact, beyond pursuing idealized theoretical analyses, there is growing interest in understanding the theoretical limits of Transformers under more realistic constraints—namely finite precision, width, and depth. Such analyses typically fall into two categories: upper bounds and lower bounds on the model’s expressive power, which we elaborate on below. For more detailed discussions and broader context, we refer to the recent survey literature [represent-survey].
@@
Upper bounds. Upper bounds typically rely on specific constructions to demonstrate that a model can represent a given function or solve a given task, i.e., that a valid solution indeed exists. Once such a solution is found, it provides an upper limit on the model size required to realize this capability [hahn2020theoretical, bhattamishra2020ability, yao2021self, wen2023transformers, liu2022transformers]. Because these results often rely on constructive proofs, the tasks or problems they address are typically case-by-case. Although self-attention networks cannot process formal languages with hierarchical structure [hahn2020theoretical, bhattamishra2020ability], such as $\rm{Dyck}_{k}$, yao2021self demonstrate that they can process $\rm{Dyck}_{k,D}$, the subset of $\rm{Dyck}_{k}$ with depth bounded by $D$. wen2023transformers show that different attention patterns can be Learned to generate bounded $\rm{Dyck}$ and interpretability via local (“myopic”) analysis can be provably misleading on Transformers. liu2022transformers prove that a shallow Transformer with $o(T)$ layers can exactly simulate any finite-state automaton processing a sequence of length $T$ and $O(\log T)$-depth Transformers always exist to simulate any automaton of length $T$. Moreover, even $O(1)$-depth solutions are surprisingly common.
@@
Lower bounds. Lower bounds show that any model capable of effectively representing a certain function or solving a certain task must have a size at least as large as some specified threshold. A useful approach is to view the model as a circuit, allowing its expressive power to be characterized using circuit complexity, which studies the minimum circuit size or depth required to solve a computational problem [merrill2023parallelism, li2024chain, hao2022formal, hao2022formal]. merrill2023parallelism show that Transformers with $O(1)$-depth and log-precision can only solve problems within the class $TC^{0}$. If the precision is further restricted to be constant, then such Transformers are limited to solving problems in $AC^{0}$ as shown by li2024chain. hao2022formal prove that generalized UHAT (unique hard attention) models can only recognize $AC^{0}$ languages while an averaging hard-attention (AHAT) model that can recognize non-$AC^{0}$ languages. Another line of work approaches the question through the lens of communication complexity. These studies recast the anchoring problem into a known communication problem and then identify the resulting communication bottlenecks imposed by model width [sanford2023representational, sanford2024transformers, peng2024limitations, chen2024theoretical, chen2024theoretical]. sanford2023representational introduce the “sparse averaging” task and show that Transformers achieve only $O(\log T)$ communication complexity, in contrast to the polynomial requirements of RNNs and feed-forward networks. Furthermore, sanford2024transformers demonstrate that log-depth Transformers can solve some basic computational tasks that cannot be efficiently handled by other neural sequence models or by sub-quadratic Transformer approximations. peng2024limitations employ communication complexity to show that a Transformer layer cannot reliably compose functions once the function domains become sufficiently large, a limitation that has been linked to the emergence of hallucinations. chen2024theoretical further provide the first unconditional lower bound for multi-layer decoder-only Transformers: for any fixed depth $L$, an $L$-layer Transformer must have polynomial width in order to perform the sequential composition of $L$ functions over $T$ tokens. In addition, by reducing the in-context learning problem to a set-disjointness task, arora2024just demonstrate that a recurrent model’s ability to recall information is sensitive to the order in which inputs are presented.

### v2

@@ 2.2.3 Memorization
Beyond the practical strategy of deduplication, a core theoretical issue in data preparation is the intrinsic mechanism of memorization (Wei et al., 2024). While often viewed as a privacy risk, memorization is deeply intertwined with the model’s learning and generalization capabilities. Research in this area has evolved from observing exact replication to analyzing complex memory representations, quantification methods, and its fundamental trade-offs with generalization.
@@
The Mechanism of Memorization. Academic discourse first challenged the traditional view that memorization is caused solely by exact sequence duplication in the training data. A foundational study introduced the concept of “Mosaic Memory” (Shilov et al., 2024). This work posits that LLM memorization is not merely verbatim recall, but a more complex process where models can patch together memories by integrating partially overlapping or similar sequences (i.e., fuzzy duplicates) from the training corpus. Building on this, other research redefined memorization from an adversarial perspective, proposing the “Adversarial Compression Ratio” (ACR) (Schwarzschild et al., 2024). The core idea is that a training sequence is considered “memorized” if it can be elicited by a prompt that is significantly shorter than the string itself. This metric provides a practical, adversarial view for assessing data usage compliance and potential privacy violations.
@@
Quantify the Influence of Memorization. Once complex memory forms were defined, the focus shifted to its quantification and prediction. Carlini et al. (2022) confirm that memorization is more prevalent than previously believed and is likely to get worse as models continue to scale, at least without active mitigation. This scaling behavior was specifically quantified for factual knowledge, with one study proposing “Scaling Laws for Fact Memorization” (Lu et al., 2024). It found that a model’s fact knowledge capacity exhibits a linear relationship with model size and a negative exponential relationship with training epochs. Beyond model scale, data-side characteristics are also a critical factor. The “Entropy-Memorization Law” (Huang et al., 2025) was proposed to investigate the inherent difficulty of memorizing data. This law reveals a linear correlation: the data’s entropy is linearly correlated with its memorization score, suggesting that simpler, lower-entropy data is more easily memorized. In terms of predictability, Biderman et al. (2023) have further shown that using a partially trained model to predict memorization is more effective than using a small model.
@@
Memorization Analysis. The ultimate goal of understanding memorization is to differentiate it from generalization and to enable effective control. A key challenge is diagnosing memorization in black-box models. The PEARL (Djiré et al., 2025) framework was introduced as a novel detection method based on a perturbation sensitivity hypothesis. This hypothesis posits that memorized content is more sensitive to input perturbations, whereas generalized knowledge remains robust. Further analysis reveals a clear trade-off between memorization and generalization across different tasks. Wang et al. (2025d) traced model capabilities back to pretraining data, finding that task dependencies vary significantly. For instance, Factual Question Answering demonstrates the strongest memorization effect, and this effect increases with model size. Conversely, tasks like machine translation and reasoning exhibit greater generalization, tending to produce novel outputs. Based on these theoretical insights, researchers have begun exploring active mitigation strategies. “Memorization Sinks” (Ghosal et al., 2025) is proposed to activate a unique set of “memorization neurons” for each sequence. This mechanism effectively isolates the memorized content, making it easier to remove without compromising general language capabilities, offering a new path to mitigate the negative impacts of memorization.
@@ 2.3.1 Synthetic Data Generation
One of the most exciting and debated frontiers is the use of synthetic data and the potential for a self-improvement loop (Villalobos et al., 2024; Long et al., 2024). Can a model generate new, high-quality data to train its next generation, thereby kicking off a cycle of recursive self-improvement? While numerous works adopt synthetic data to improve model training (Patel et al., 2024; Gilardi et al., 2023; Xu et al., 2023; Gandhi et al., 2024; Liu et al., 2024b), this idea faces significant theoretical hurdles. A key open question is whether such a process would lead to genuine capability gains or result in model collapse, a degenerative process where the model overfits to its own idiosyncrasies, leading to a gradual loss of diversity and accuracy. Developing a theoretical framework to understand and control the dynamics of this loop is a critical area of research.
@@
Recent researches have started to establish theoretical frameworks for the utility of synthetic data. Gan and Liu (2025) propose a “reverse-bottleneck” framework, which posits that a post-trained model’s generalization error upper bound is negatively correlated with the “information gain” obtained from the generative model. This suggests that so long as the generative model provides sufficient new information, generalization can in principle be improved. Beyond simple augmentation, the nature of the synthetic data is also being explored. For instance, in the domain of mathematical reasoning, Setlur et al. (2024) find that while fine-tuning on correct synthetic answers offers modest gains, using reinforcement learning on the model’s incorrect responses can be twice as sample-efficient. This method helps the model identify and unlearn “spurious correlations” (i.e., incorrect intermediate steps that happen to lead to a correct final answer), ultimately scaling the synthetic dataset’s efficiency by eight-fold compared to standard positive-only finetuning.
@@
The primary theoretical hurdle to recursive self-improvement is “Model Collapse”. Shumailov et al. (2023) provide a foundational study on this phenomenon, positing that training on generated data leads to an irreversible degenerative process. Beyond this, other works have highlighted the limitations of synthetic data in capturing human nuance. Li et al. (2023d) find that the performance gap between real and synthetic data is smallest for low-subjectivity tasks (like news classification) but much larger for high-subjectivity tasks (like humor or sarcasm detection). This suggests LLMs struggle to generate data with sufficient diversity to capture the complexity of subjective language. This is empirically supported by Møller et al. (2024), which finds that models trained on human-labeled data consistently exhibited superior or comparable performance to those trained on synthetically augmented data.
@@
In response to the threat of model collapse, various mitigation strategies have emerged. Multiple studies have empirically and theoretically demonstrated that the training workflow is the critical factor (Gerstgrasser et al., 2024; Kazdan et al., 2024). A “replace” workflow, which discards old data and trains new models only on synthetic data, does lead to collapse. However, an “accumulate” workflow, where synthetic data is added alongside the original real data, consistently avoids model collapse and keeps models stable. Seddik et al. (2024) provide a quantitative estimate, concluding that to maintain stability, the amount of synthetic data used must be considerably smaller than the amount of real data in the training mix. This body of work suggests that the value of synthetic data is highly context-dependent: it can improve performance when real data is scarce but may harm it when real data is plentiful.
@@ 2.3.2 Data Contamination
Data contamination, the inadvertent inclusion of benchmark evaluation samples within the pre-training corpus, poses another critical open challenge in the data preparation stage. This issue fundamentally undermines the validity of model evaluations, making it difficult to discern true generalization capabilities from mere memorization of seen answers (Deng et al., 2024a; Cheng et al., 2025b; Xu et al., 2024a). The theoretical and empirical investigation of contamination can be broadly categorized by its severe impacts and the methods for its detection and mitigation.
@@
Data contamination is a direct threat to reliable model assessment. Studies demonstrate that contamination can drastically skew evaluation scores. Kocyigit et al. (2025) find that contamination in a machine translation task could severely inflate model capability. This work also revealed that larger models exhibit higher sensitivity to contamination, not greater robustness. Li and Flanigan (2024) further enrich the evaluation, they find that a model’s superior performance in apparent zero- or few-shot settings may not stem from genuine generalization but from its exposure to task-related samples during pre-training. The impact on complex reasoning evaluation is particularly stark. Huang et al. (2024b) test models on novel competition problems released after their training data cut-off. They find a “cliff like decline” in GPT-4’s performance on medium-to-hard problems, strongly suggesting that its high performance on older benchmarks was reliant on memorization rather than genuine algorithmic reasoning. However, contamination is not limited to harmful verbatim copies. Palavalli et al. (2024) establish through experiments that “noisy” or approximate forms of contamination (e.g., masking, augmenting, or noising test examples) can boost performance almost as much as seeing clean, in-domain data. Beyond evaluation, the memorization of contaminated data, especially sensitive information, creates significant privacy vulnerabilities. Zhu et al. (2024a) provide a systematic benchmark for assessing these privacy leakage risks, which are exacerbated during model adaptation and fine-tuning.
@@
In response to these severe impacts, researchers have developed various methods for detection, though effective mitigation remains a significant open problem. Choi et al. (2025) propose the KDS framework. Instead of simple text matching, it quantifies contamination by measuring the change in the similarity structure of sample embeddings in the model’s representation space before and after fine-tuning. Deng et al. (2024b) introduce an innovative detection method called TS-Guessing. This protocol masks an incorrect answer in a multiple-choice question and prompts the model to fill in the blank. Commercial LLMs were able to “guess” the exact missing wrong option with high accuracy, strongly implying they had memorized the full question format. Other approaches use targeted queries to excavate a model’s memory. Chang et al. (2023) use “name cloze” queries to identify a wide range of memorized copyrighted books, which in turn contaminated downstream evaluation tasks. Similarly, Dodge et al. (2021) confirm that the C4 dataset (Raffel et al., 2020a) contains contaminated examples from NLP benchmarks. Contamination is not just a pre-training issue. Tao et al. (2025) address the challenge of detection after RLHF, where optimization erases traditional likelihood signals. The proposed “Self-Critique” method probes for “policy collapse” by comparing the token-level entropy sequences of an initial response and a second, alternative critique response, where high similarity indicates memorization.
@@ 3.1 Fundamental Problems
(1) How to theoretically evaluate the “power" of a model architecture? This problem focuses on the rigorous analysis of a model’s intrinsic properties, specifically its representability: the capacity to solve or approximate given classes of functions. The challenge is to determine the theoretical limits of architectures under realistic constraints such as finite precision, width, and depth. This involves establishing upper bounds on the model size required to realize specific capabilities and lower bounds that characterize the minimum circuit or communication complexity needed to solve computational tasks.
@@
Of course, discussions on these questions often overlap. For example, according to the “no free lunch" principle, there is usually a trade-off between model performance and computational cost, that is, efficient models may come with potential limitations in representational power. Moreover, the theoretical understanding of a model’s architecture is closely related to the observable characteristics it exhibits. We present a visualization of the topics in this stage in fig.˜3. In what follows, we start with the core theories and methods relevant to model design.
@@ 3.2.1 Representability of Models
The community has renewed interest in the expressive capacity of Transformers, especially in terms of universal approximation (Yun et al., 2019; Jiang and Li, 2024; Kajitsuka and Sato, 2023). Yun et al. (2019) show that for any sequence-to-sequence function, there exists a Transformer that can approximate it, where the number of layers scales exponentially in model dimension $d$ or input sequence length $T$ and the size of each layer is independent of $d$ and $T$. Jiang and Li (2024) derive explicit Jackson-type approximation bounds for Transformers by introducing new complexity measures to construct appropriate approximation spaces and their results show that Transformers approximate efficiently when the temporal dependencies of the target function exhibit a low-rank structure. Kajitsuka and Sato (2023) demonstrate that once the feed-forward layer quantizes continuous inputs, even a one-layer, single-head self-attention module becomes a universal approximator for continuous permutation equivariant functions on a compact domain. More recently, Su et al. (2026); Su and Liu (2026) also extend the representability analysis towards more complex mixture-of-experts (MoE) architectures by introducing tropical geometry.
@@
In addition, there are also many works that study representability through Turing completeness, asking whether a model can effectively simulate each step of a Turing machine—and thus inherit the full computational power of Turing-computable problems (Dehghani et al., 2018; Pérez et al., 2021; Wei et al., 2022a). Pérez et al. (2021) prove that Transformers are indeed Turing complete under the assumption of infinite precision. Dehghani et al. (2018) show that standard finite-precision Transformers are not Turing complete and the proposed Universal Transformer which combines parallel self-attention with recurrence can overcome this limitation. Wei et al. (2022a) demonstrate that Transformers can statistically meaningfully approximate Turing machines running in time $O(T)$, with sample complexity polynomial in the alphabet size, state-space size, and $\log(T)$. In fact, beyond pursuing idealized theoretical analyses, there is growing interest in understanding the theoretical limits of Transformers under more realistic constraints—namely finite precision, width, and depth. Such analyses typically fall into two categories: upper bounds and lower bounds on the model’s expressive power, which we elaborate on below. For more detailed discussions and broader context, we refer to the recent survey literature (Strobl et al., 2024).
@@
Upper bounds. Upper bounds typically rely on specific constructions to demonstrate that a model can represent a given function or solve a given task, i.e., that a valid solution indeed exists. Once such a solution is found, it provides an upper limit on the model size required to realize this capability (Hahn, 2020; Bhattamishra et al., 2020; Yao et al., 2021; Wen et al., 2023; Liu et al., 2022). Because these results often rely on constructive proofs, the tasks or problems they address are typically case-by-case. Although self-attention networks cannot process formal languages with hierarchical structure (Hahn, 2020; Bhattamishra et al., 2020), such as $\rm{Dyck}_{k}$, Yao et al. (2021) demonstrate that they can process $\rm{Dyck}_{k,D}$, the subset of $\rm{Dyck}_{k}$ with depth bounded by $D$. Wen et al. (2023) show that different attention patterns can be Learned to generate bounded $\rm{Dyck}$ and interpretability via local (“myopic") analysis can be provably misleading on Transformers. Liu et al. (2022) prove that a shallow Transformer with $o(T)$ layers can exactly simulate any finite-state automaton processing a sequence of length $T$ and $O(\log T)$-depth Transformers always exist to simulate any automaton of length $T$. Moreover, even $O(1)$-depth solutions are surprisingly common.
@@
Lower bounds. Lower bounds show that any model capable of effectively representing a certain function or solving a certain task must have a size at least as large as some specified threshold. A useful approach is to view the model as a circuit, allowing its expressive power to be characterized using circuit complexity, which studies the minimum circuit size or depth required to solve a computational problem (Merrill and Sabharwal, 2023b; Li et al., 2024b; Hao et al., 2022). Merrill and Sabharwal (2023b) show that Transformers with $O(1)$-depth and log-precision can only solve problems within the class $TC^{0}$. If the precision is further restricted to be constant, then such Transformers are limited to solving problems in $AC^{0}$ as shown by Li et al. (2024b). Hao et al. (2022) prove that generalized UHAT (unique hard attention) models can only recognize $AC^{0}$ languages while an averaging hard-attention (AHAT) model that can recognize non-$AC^{0}$ languages. Another line of work approaches the question through the lens of communication complexity. These studies recast the anchoring problem into a known communication problem and then identify the resulting communication bottlenecks imposed by model width (Sanford et al., 2023; 2024; Peng et al., 2024; Chen et al., 2024b; b). Sanford et al. (2023) introduce the “sparse averaging” task and show that Transformers achieve only $O(\log T)$ communication complexity, in contrast to the polynomial requirements of RNNs and feed-forward networks. Furthermore, Sanford et al. (2024) demonstrate that log-depth Transformers can solve some basic computational tasks that cannot be efficiently handled by other neural sequence models or by sub-quadratic Transformer approximations. Peng et al. (2024) employ communication complexity to show that a Transformer layer cannot reliably compose functions once the function domains become sufficiently large, a limitation that has been linked to the emergence of hallucinations. Chen et al. (2024b) further provide the first unconditional lower bound for multi-layer decoder-only Transformers: for any fixed depth $L$, an $L$-layer Transformer must have polynomial width in order to perform the sequential composition of $L$ functions over $T$ tokens. In addition, by reducing the in-context learning problem to a set-disjointness task, Arora et al. (2024) demonstrate that a recurrent model’s ability to recall information is sensitive to the order in which inputs are presented.